4Er+3O2->2Er2O3



how many gram of Er must be reacted with O2 to produce 28.78gEr2O3?

1
Expert's answer
2021-11-12T13:48:03-0500

M(Er2O3) = 382.56 g/mol

n(Er2O3) "= \frac{28.78}{382.56} = 0.075 \;mol"

According to the reaction:

n(Er) = 2n(Er2O3) "= 2 \times 0.075 = 0.150 \;mol"

M(Er) = 167.26 g/mol

m(Er) "= 0.150 \times 167.26 = 25.16 \;g"

Answer: 25.16 g